OpenStudy (winsteria):

MEDAL AND FAN What is the ratio of x to y? 3/2 =2x/2x+y A. -3/2 B. -2/3 C. 2/3 D. 3/2

OpenStudy (winsteria):

@AlexandervonHumboldt2 @iGreen

OpenStudy (igreen):

Wait..is this it? \(\dfrac{3}{2} = \dfrac{2x}{2x + y}\)

OpenStudy (winsteria):

Yes

OpenStudy (igreen):

Cross multiply

OpenStudy (winsteria):

cross multiply what though?

OpenStudy (winsteria):

@iGreen

OpenStudy (igreen):

3 * (2x + y) = 2 * 2x

OpenStudy (winsteria):

6 and 4?

OpenStudy (igreen):

Distribute 3 into the parenthesis. 3 * 2x = ? 3 * y = ?

OpenStudy (winsteria):

6x?

OpenStudy (igreen):

Yep, what about 3 * y

OpenStudy (winsteria):

I don't know what y is though

OpenStudy (winsteria):

@iGreen

OpenStudy (winsteria):

@TheSmartOne

OpenStudy (winsteria):

@KendrickLamar2014

OpenStudy (igreen):

3 * y = 3y

OpenStudy (igreen):

So we have: 6x + 3y = 2 * 2x What's 2 * 2x?

OpenStudy (winsteria):

4x

OpenStudy (winsteria):

what's next?

OpenStudy (winsteria):

@iGreen

OpenStudy (igreen):

We have: 6x + 3y = 4x Subtract 6x to both sides, what's 4x - 6x?

OpenStudy (winsteria):

2x

OpenStudy (winsteria):

@iGreen

OpenStudy (igreen):

Actually that's negative, which gives us: 3y = -2x Divide -2 to both sides: \(-\dfrac{3}{2}y = x\) Divide 'y' to both sides: \(-\dfrac{3}{2} = \dfrac{x}{y}\) So -3/2 is your answer.

OpenStudy (winsteria):

Thank you